# Enhanced Phase Estimation via Photon-Added Two-Mode Squeezed States and Kerr Nonlinearity

### v1

@@ Abstract
Quantum metrology harnesses quantum resources to achieve measurement precision beyond classical limits. This work investigates a Mach-Zehnder interferometer incorporating a Kerr nonlinear phase shifter, using photon-added two-mode squeezed coherent states generated via four-wave mixing as input. This study demonstrates that increasing both the photon-addition order and input resource strength systematically enhances phase sensitivity, quantum Fisher information, and the quantum Cramér-Rao bound. The system not only surpasses the standard quantum limit but also approaches or exceeds the Heisenberg limit with linear phase shifts, while Kerr nonlinearity enables overcoming the super-Heisenberg limit. The proposed scheme exhibits enhanced robustness against photon loss, providing a promising approach for high-precision quantum metrology applications.
@@ I Introduction
Precision measurement constitutes the foundation of scientific and technological advancement, with its fundamental limits defining our capacity to understand and manipulate physical phenomena. From traditional mechanical manufacturing to modern lithography techniques, and from gravitational wave detection to biomedical sensing, technological progress consistently relies on advances in measurement precision. Within classical frameworks, however, measurement accuracy is fundamentally constrained by the standard quantum limit (SQL), where measurement uncertainty scales inversely with the square root of the average photon number $\bar{N}$ within the interferometer 1 ; 2 ; 3 .
@@
Quantum mechanics provides transformative approaches for overcoming these limitations. Quantum metrology exploits distinctive quantum resources—including superposition, entanglement, and non-classical correlations—to surpass the SQL barrier and achieve ultra-high precision measurements approaching or exceeding the Heisenberg limit (HL), scaling as $1/\bar{N}$ 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 ; 11 . This field possesses profound physical implications and demonstrates substantial potential in applications ranging from gravitational wave detection and atomic clocks to magnetometry, super-resolution imaging, and measurements of fundamental physical constants.
@@
A principal strategy for quantum enhancement involves injecting non-classical states into interferometric devices rather than conventional coherent states. Squeezed states and entangled states generated through parametric amplification processes have proven effective for surpassing the SQL. However, preparing ideal entangled states (such as NOON states) remains challenging due to their extreme susceptibility to environmental noise and photon loss, significantly limiting practical implementation 11 ; 12 ; 13 ; 14 ; 15 ; 16 .
@@
Recent research has extensively explored non-Gaussian operations, particularly photon addition and subtraction, which can substantially enhance quantum states’ non-classical properties and potentially improve measurement robustness. Among these, photon addition operations offer distinct advantages in quantum state preparation and in enhancing both non-classical characteristics and measurement precision 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 . Concurrently, interferometer architectures have been continuously refined. For instance, incorporating Kerr nonlinear media to replace conventional linear phase shifters can theoretically achieve accuracy beyond the sub-HL ($1/\bar{N}^{3/2}$), potentially reaching the super-Heisenberg limit (SHL), i.e., $1/\bar{N}^{2}$ 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 .
@@
Building upon this foundation, this work develops a quantum precision measurement scheme combining high precision with enhanced robustness. We generate two-mode squeezed coherent states (TMSCS) via four-wave mixing (FWM) and apply photon addition operations to create photon-added TMSCS (PA-TMSCS) as the interferometer input resource. Furthermore, we replace the linear phase shifter in a conventional Mach-Zehnder interferometer (MZI) with a Kerr nonlinear phase shifter, constructing a Kerr nonlinear MZI (KMZI). This study systematically analyzes the scheme’s performance under both ideal conditions and in the presence of photon loss. Through examining phase sensitivity via intensity difference detection and quantum Fisher information (QFI), and comparing linear and nonlinear phase shift scenarios, we verify the synergistic advantages of photon addition operations and Kerr nonlinear effects in enhancing measurement accuracy and system resilience. This work provides valuable theoretical insights for developing practical quantum-enhanced measurement technologies.
@@
The paper is organized as follows: Section II introduces the phase estimation model based on a KMZI with PA-TMSCS input. Section III investigates phase sensitivity of intensity difference detection for both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, including photon loss effects. Section IV examines the QFI for these cases under ideal and lossy conditions. Conclusions are presented in the final section.
@@ II Phase Estimation Model Based on a KMZI with PA-TMSCS Input
Figure 1: Schematic diagram of phase estimation using PA-TMSCS inputs in a KMZI, where photon addition operations on TMSCS in both $a$-mode and $b$-mode are considered. Photon loss within the KMZI is simulated by a virtual beam splitter, and intensity difference detection is implemented at the output port.
This section presents the phase estimation model employing PA-TMSCS inputs in a KMZI. As illustrated in Fig. 1, TMSCS can be generated by applying an FWM process induced by a pump field to coherent states. The FWM process is equivalent to a two-mode squeezing operation described by the operator $S_{2}\left(r\right)=\exp\left[\xi^{\ast}ab-\xi a^{\dagger}b^{\dagger}\right]$, where $\xi=re^{i\phi_{1}}$, with $r$ and $\phi_{1}$ representing the squeezing parameter and phase, respectively. To simplify experimental procedures and avoid complicated phase-matching conditions, only the $a$-mode input is prepared as a coherent state $\left|\alpha\right\rangle_{a}$, while the $b$-mode input remains in the vacuum state. The coherent state satisfies $a\left|\alpha\right\rangle_{a}=\alpha\left|\alpha\right\rangle_{a}$, where $\alpha=\left|\alpha\right|e^{i\theta}$ ($\left|\alpha\right|$ and $\theta$ denote amplitude and phase, respectively). For simplicity and to satisfy phase matching conditions, we set $\theta=0$ ($\alpha=\left|\alpha\right|$) and $\phi_{1}=\pi$, which proves advantageous for enhanced phase estimation 24 ; 42 .
To further optimize input states for improved measurement accuracy, we implement non-Gaussian photon addition operations on TMSCS. The resulting PA-TMSCS serves as the input state and can be expressed as
$\left|\text{in}\right\rangle=\frac{1}{\sqrt{N_{m,n}}}a^{\dagger m}b^{\dagger n}S_{2}\left(r\right)\left|\alpha\right\rangle_{a}\left|0\right\rangle_{b},$ (1)
where $a^{\dagger m}$ and $b^{\dagger n}$ represent photon addition operations on the $a$-mode and $b$-mode, respectively, and $N_{m,n}$ is the normalization coefficient. To facilitate calculations, the photon addition operators in Eq. (1) can be converted to partial derivative form:
$\left|\text{in}\right\rangle$ $=$ $\frac{1}{\sqrt{N_{m,n}}}\frac{\partial^{m+n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}}\exp\left[a^{\dagger}t_{1}\right]$ (2)
$\times\exp\left[b^{\dagger}\tau_{1}\right]S_{2}\left(r\right)\left|\alpha\right\rangle_{a}\left|0\right\rangle_{b}|_{t_{1}=\tau_{1}=0}.$
Based on this representation and quantum state normalization, the expression for $N_{m,n}$ can be derived through relevant operator transformation relations:
$\begin{array}[]{c}N_{m,n}=\Gamma_{0}\{\exp[((t_{1}t_{2}+\tau_{1}\tau_{2})\cosh r\\ +(t_{1}\tau_{1}+t_{2}\tau_{2})\sinh r)\cosh r]\\ \times\exp[((t_{1}+t_{2})\cosh r\\ +(\tau_{1}+\tau_{2})\sinh r)\alpha]\},\end{array}$ (3)
@@
$\Gamma_{0}\left\{\cdot\right\}=\frac{\partial^{2m+2n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}}\left.\left\{\cdot\right\}\right|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=0}.$ (4)
Conventional MZIs constitute linear measuring instruments comprising two 50:50 beam splitters (BS1 and BS2) and a linear phase shifter. Following Yurke et al. 43 , using angular momentum operators in the Schwinger representation, the equivalent operators for BS1 and BS2 are $B_{1}=\exp\left[-i\pi(a^{\dagger}b+ab^{\dagger})/4\right]$ and $B_{2}=\exp\left[i\pi(a^{\dagger}b+ab^{\dagger})/4\right]$, respectively. These operators satisfy the transformation relations:
$B_{1}^{\dagger}\binom{a}{b}B_{1}=\frac{\sqrt{2}}{2}\left(\begin{array}[]{cc}1&-i\\ -i&1\end{array}\right)\binom{a}{b}$, (5)
@@
$B_{2}^{\dagger}\binom{a}{b}B_{2}=\frac{\sqrt{2}}{2}\left(\begin{array}[]{cc}1&i\\ i&1\end{array}\right)\binom{a}{b}$. (6)
For our investigation, we replace the linear phase shifter in the conventional MZI with a Kerr nonlinear phase shifter to construct an improved KMZI. The phase shifter’s equivalent operator is $U\left(\phi,k\right)=\exp[i\phi(a^{\dagger}a)^{k}]$, where $k=1$ corresponds to linear phase shift and $k=2$ to Kerr nonlinear phase shift. Here, $\phi$ represents the phase shift to be measured. The transformation relations for linear and Kerr nonlinear phase shifts are respectively given by:
$U^{\dagger}\left(\phi,1\right)aU\left(\phi,1\right)=e^{i\phi}a,$ (7)
@@
$U^{\dagger}\left(\phi,2\right)aU\left(\phi,2\right)=e^{i\phi}e^{i\phi\left(2a^{\dagger}a\right)}a.$ (8)
The transformation relation for $U\left(\phi,2\right)$ in Eq. (8) can be derived using methods from Ref. 40 based on Fock state representation, with detailed steps provided in Appendix A.
@@
$B_{f}^{\dagger}\left(\begin{array}[]{c}a\\ a_{f}\end{array}\right)B_{f}=\left(\begin{array}[]{cc}\sqrt{1-l}&\sqrt{l}\\ -\sqrt{l}&\sqrt{1-l}\end{array}\right)\left(\begin{array}[]{c}a\\ a_{f}\end{array}\right),$ (9)
@@
where $a_{f}$ is the photon annihilation operator for the auxiliary mode $f$ containing vacuum noise $\left|0\right\rangle_{f}$, while $B_{f}$ and $l$ represent the equivalent operator and reflectivity of the virtual beam splitter, respectively. The parameter $l$ corresponds to the photon loss rate, satisfying $0\leq l\leq 1$. Higher $l$ values indicate increased loss, with $l=0$ representing the lossless case and $l=1$ corresponding to complete absorption.
@@
Within this KMZI framework, intensity difference detection is implemented by placing photon counters at the output ports. Since intensity difference detection is relatively straightforward to implement experimentally, it finds widespread use for phase sensitivity measurements in MZIs.
III Phase Sensitivity of Intensity Difference Detection
@@
The phase sensitivity for intensity difference detection can be expressed using the error propagation formula:
@@
$\Delta\phi=\frac{\sqrt{\left\langle I_{D}^{2}\right\rangle-\left\langle I_{D}\right\rangle^{2}}}{\left|\frac{\partial\left\langle I_{D}\right\rangle}{\partial\phi}\right|},$ (10)
@@
where $I_{D}=a^{\dagger}a-b^{\dagger}b$, and $\left\langle\cdot\right\rangle=\left\langle\text{out}\right|\cdot\left|\text{out}\right\rangle$ denotes the expectation value with respect to the output state. From the phase estimation model description, the output state can be expressed as:
$\left|\text{out}\right\rangle=B_{2}U\left(\phi,k\right)B_{f}B_{1}\left|\text{in}\right\rangle\left|0\right\rangle_{f}.$ (11)
@@
We focus on phase sensitivity for both $k=1$ and $k=2$ cases. For brevity, explicit expressions for phase sensitivity are omitted here. Substituting the expectation values $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$ into Eq. (10) yields the phase sensitivity. The specific expressions for $\left\langle I_{D}\right\rangle$ and $\left\langle I_{D}^{2}\right\rangle$ under $k=1$ and $k=2$ conditions are provided in Appendices B and C, respectively.
Refer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 2: Phase sensitivity $\Delta\phi$ versus phase shift $\phi$ for squeezing parameter $r=0.9$ and coherent amplitude $\alpha=1$.
@@
Fig. 2 illustrates the variation of phase sensitivity $\Delta\phi$ with phase shift $\phi$ for $k=1$ and $k=2$ cases, where Figs. 2(a) and 2(b) depict linear phase shift ($k=1$), while Figs. 2(c) and 2(d) represent Kerr nonlinear phase shift ($k=2$). According to our computational expressions and the behavior of $\Delta\phi$ in Fig. 2(a), the variation period is $\pi$. Comparing Figs. 2(a) and 2(c), curves incorporating photon addition operations show significant improvement over the black solid line ($m=n=0$). Thus, for schemes employing PA-TMSCS input states, photon addition effectively enhances phase sensitivity. Notably, the phase sensitivity variation range in Fig. 2(c) is substantially reduced compared to Fig. 2(a) due to the Kerr nonlinear effect at $k=2$ (hence Fig. 2(c) is not plotted over a full period), suggesting that Kerr nonlinear phase shifting can significantly improve phase sensitivity within certain phase shift ranges.
@@
To further investigate effective photon addition schemes, we analyze phase sensitivity characteristics under narrowed plotting ranges, as shown in Figs. 2(b) and 2(d). Phase sensitivity improves with increasing photon addition order $m+n$, and near the optimal phase sensitivity point, symmetric photon addition ($m=n$) yields the best performance for a given $m+n$. This indicates that performing symmetric photon addition operations in both $a$- and $b$-modes is more conducive to enhancing measurement accuracy. Consequently, we focus on the $m=n$ scenario for photon addition operations throughout this paper.
Refer to captionRefer to caption
@@
Figure 3: For fixed phase shift $\phi=3.12$: (a) phase sensitivity $\Delta\phi$ versus squeezing parameter $r$ with coherent amplitude $\alpha=1$; (b) $\Delta\phi$ versus $\alpha$ with $r=0.9$.
@@
Fig. 3 demonstrates that phase sensitivity $\Delta\phi$ improves with increasing squeezing parameter $r$, coherent amplitude $\alpha$, and photon addition order. Additionally, comparing solid and dotted lines clearly shows that Kerr nonlinear phase shift provides improved phase sensitivity compared to linear phase shift.
@@
In practical scenarios, photon loss constitutes an inevitable factor. We subsequently analyze its impact on phase sensitivity and compare achievable measurement precision limits. The measurement precision limit can be defined through the average photon number within the interferometer $\bar{N}$, expressed as:
$\bar{N}$ $=$ $\left.{}_{\text{int}}\left\langle\psi\right|a^{\dagger}a+b^{\dagger}b\left|\psi\right\rangle_{\text{int}}\right$. (12)
$=$ $D_{1}\left(1,0,1,0\right)+D_{1}\left(0,1,0,1\right),$
@@
where $\left|\psi\right\rangle_{\text{int}}=U\left(\phi,k\right)B_{f}B_{1}\left|\text{in}\right\rangle\left|0\right\rangle_{f}$, and specific expressions for $D_{1}\left(1,0,1,0\right)$ and $D_{1}\left(0,1,0,1\right)$ are given in Eq. (B5) of Appendix B.
Refer to captionRefer to captionRefer to caption
@@
Figure 4: Linear phase shift case: phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL and HL are shown for comparison.
Refer to captionRefer to captionRefer to caption
@@
Figure 5: Kerr nonlinear phase shift case: phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL, HL, sub-HL, and SHL are shown for comparison.

### v2

@@ Abstract
Quantum metrology employs quantum resources to achieve measurement precision beyond classical limits. This work investigates a Mach–Zehnder interferometer incorporating a Kerr nonlinear phase shifter, with photon-added two-mode squeezed coherent states generated via four-wave mixing as input. We demonstrate that increasing both the photon-addition order and the input resource strength systematically enhances phase sensitivity, quantum Fisher information, and the corresponding quantum Cramér–Rao bound. The proposed system not only surpasses the standard quantum limit but also approaches or exceeds the Heisenberg limit for linear phase shifts, while Kerr nonlinearity enables surpassing the super-Heisenberg limit. Furthermore, the scheme exhibits enhanced robustness against photon loss, providing a promising pathway toward practical high-precision quantum metrology applications.
@@ I Introduction
Precision measurement constitutes the cornerstone of scientific and technological advancement, with its fundamental limits dictating our ability to probe and manipulate physical phenomena. From traditional mechanical manufacturing to modern lithography, and from gravitational-wave detection to biomedical sensing, progress consistently relies on advancing measurement precision. Within classical frameworks, measurement accuracy is fundamentally constrained by the standard quantum limit (SQL), where measurement uncertainty scales as $1/\sqrt{\bar{N}}$ with the average photon number $\bar{N}$ inside the interferometer 1 ; 2 ; 3 .
@@
Quantum mechanics offers transformative strategies to overcome these limitations. Quantum metrology exploits distinct quantum resources—including superposition, entanglement, and nonclassical correlations—to surpass the SQL and achieve ultra-high precision approaching or exceeding the Heisenberg limit (HL), which scales as $1/\bar{N}$ 4 ; 5 ; 6 ; 7 ; 8 ; 9 ; 10 ; 11 . This field carries profound physical implications and substantial potential for applications ranging from gravitational-wave detection and atomic clocks to magnetometry, super-resolution imaging, and measurements of fundamental physical constants.
@@
A primary strategy for quantum enhancement involves injecting nonclassical states into interferometric devices rather than conventional coherent states. Squeezed states and entangled states generated via parametric amplification processes have proven effective in surpassing the SQL. However, preparing ideal entangled states (e.g., NOON states) remains challenging due to their extreme susceptibility to environmental noise and photon loss, which significantly limits practical implementation 11 ; 12 ; 13 ; 14 ; 15 ; 16 .
@@
Recent research has extensively explored non-Gaussian operations, particularly photon addition and subtraction, which can substantially enhance the nonclassical properties of quantum states and potentially improve measurement robustness. Among these, photon-addition operations offer distinct advantages in quantum-state preparation and in enhancing both nonclassical characteristics and measurement precision 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 . Concurrently, interferometer architectures have been continuously refined. For instance, incorporating Kerr nonlinear media to replace conventional linear phase shifters can theoretically achieve accuracies beyond the sub-HL ($1/\bar{N}^{3/2}$), potentially reaching the super-Heisenberg limit (SHL), i.e., $1/\bar{N}^{2}$ 33 ; 34 ; 35 ; 36 ; 37 ; 38 ; 39 ; 40 ; 41 .
@@
Building upon this foundation, this work develops a quantum precision measurement scheme that combines high precision with enhanced robustness. We generate two-mode squeezed coherent states (TMSCS) via four-wave mixing (FWM) and apply photon-addition operations to create photon-added TMSCS (PA-TMSCS) as the interferometer input resource. Furthermore, we replace the linear phase shifter in a conventional Mach–Zehnder interferometer (MZI) with a Kerr nonlinear phase shifter, constructing a Kerr nonlinear MZI (KMZI). This study systematically analyzes the scheme’s performance under both ideal conditions and in the presence of photon loss. By examining phase sensitivity via intensity-difference detection and quantum Fisher information (QFI), and comparing linear and nonlinear phase-shift scenarios, we verify the synergistic advantages of photon-addition operations and Kerr nonlinear effects in enhancing measurement accuracy and system resilience. This work provides valuable theoretical insights for developing practical quantum-enhanced measurement technologies.
@@
The paper is organized as follows: Section II introduces the phase estimation model based on a KMZI with PA-TMSCS input. Section III investigates the phase sensitivity of intensity-difference detection for both linear ($k=1$) and Kerr nonlinear ($k=2$) phase shifts, including the effects of photon loss. Section IV examines the QFI for these cases under ideal and lossy conditions. Conclusions are presented in the final section.
@@ II Phase Estimation Model Based on a KMZI with PA-TMSCS Input
Figure 1: Schematic of phase estimation using PA-TMSCS inputs in a KMZI. Photon-addition operations are applied to both the $a$-mode and $b$-mode of the TMSCS. Photon loss inside the KMZI is simulated by a virtual beam splitter, and intensity-difference detection is implemented at the output port.
This section presents the phase estimation model employing PA-TMSCS inputs in a KMZI. As illustrated in Fig. 1, TMSCS can be generated by applying an FWM process induced by a pump field to coherent states. The FWM process is equivalent to a two-mode squeezing operation described by the operator $S_{2}(r)=\exp[\xi^{\ast}ab-\xi a^{\dagger}b^{\dagger}]$, where $\xi=re^{i\phi_{1}}$, with $r$ and $\phi_{1}$ representing the squeezing parameter and phase, respectively. To simplify experimental procedures and avoid complicated phase-matching conditions, only the $a$-mode input is prepared as a coherent state $\lvert\alpha\rangle_{a}$, while the $b$-mode input remains in the vacuum state $\lvert 0\rangle_{b}$. The coherent state satisfies $a\lvert\alpha\rangle_{a}=\alpha\lvert\alpha\rangle_{a}$, where $\alpha=\lvert\alpha\rvert e^{i\theta}$ ($\lvert\alpha\rvert$ and $\theta$ denote amplitude and phase, respectively). For simplicity and to satisfy phase-matching conditions, we set $\theta=0$ ($\alpha=\lvert\alpha\rvert$) and $\phi_{1}=\pi$, which proves advantageous for enhanced phase estimation 24 ; 42 .
To further optimize the input states for improved measurement accuracy, we implement non-Gaussian photon-addition operations on the TMSCS. The resulting PA-TMSCS serves as the input state and can be expressed as
$\lvert\text{in}\rangle=\frac{1}{\sqrt{N_{m,n}}}a^{\dagger m}b^{\dagger n}S_{2}(r)\lvert\alpha\rangle_{a}\lvert 0\rangle_{b},$ (1)
where $a^{\dagger m}$ and $b^{\dagger n}$ represent photon-addition operations on the $a$-mode and $b$-mode, respectively, and $N_{m,n}$ is the normalization coefficient. To facilitate calculations, the photon-addition operators in Eq. (1) can be converted into partial-derivative form:
$\lvert\text{in}\rangle$ $=$ $\frac{1}{\sqrt{N_{m,n}}}\frac{\partial^{m+n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}}\exp[a^{\dagger}t_{1}]$ (2)
$\times\exp[b^{\dagger}\tau_{1}]S_{2}(r)\lvert\alpha\rangle_{a}\lvert 0\rangle_{b}\Big|_{t_{1}=\tau_{1}=0}.$
Based on this representation and quantum-state normalization, the expression for $N_{m,n}$ can be derived through relevant operator transformation relations:
$N_{m,n}=\Gamma_{0}\Bigl\{$ $\exp\bigl[((t_{1}t_{2}+\tau_{1}\tau_{2})\cosh r$ (3)
$+(t_{1}\tau_{1}+t_{2}\tau_{2})\sinh r)\cosh r\bigr]$
$\times\exp\bigl[((t_{1}+t_{2})\cosh r$
$+(\tau_{1}+\tau_{2})\sinh r)\alpha\bigr]\Bigr\},$
@@
$\Gamma_{0}\{\cdot\}=\frac{\partial^{2m+2n}}{\partial t_{1}^{m}\partial\tau_{1}^{n}\partial t_{2}^{m}\partial\tau_{2}^{n}}\{\cdot\}\Big|_{t_{1}=\tau_{1}=t_{2}=\tau_{2}=0}.$ (4)
Conventional MZIs constitute linear measuring instruments comprising two 50:50 beam splitters (BS1 and BS2) and a linear phase shifter. Following Yurke et al. 43 , using angular momentum operators in the Schwinger representation, the equivalent operators for BS1 and BS2 are $B_{1}=\exp[-i\pi(a^{\dagger}b+ab^{\dagger})/4]$ and $B_{2}=\exp[i\pi(a^{\dagger}b+ab^{\dagger})/4]$, respectively. These operators satisfy the transformation relations:
$B_{1}^{\dagger}\binom{a}{b}B_{1}=\frac{\sqrt{2}}{2}\begin{pmatrix}1&-i\\ -i&1\end{pmatrix}\binom{a}{b}$, (5)
@@
$B_{2}^{\dagger}\binom{a}{b}B_{2}=\frac{\sqrt{2}}{2}\begin{pmatrix}1&i\\ i&1\end{pmatrix}\binom{a}{b}$. (6)
For our investigation, we replace the linear phase shifter in the conventional MZI with a Kerr nonlinear phase shifter to construct an improved KMZI. The phase shifter’s equivalent operator is $U(\phi,k)=\exp[i\phi(a^{\dagger}a)^{k}]$, where $k=1$ corresponds to a linear phase shift and $k=2$ to a Kerr nonlinear phase shift. Here, $\phi$ represents the phase shift to be measured. The transformation relations for linear and Kerr nonlinear phase shifts are respectively given by:
$U^{\dagger}(\phi,1)aU(\phi,1)=e^{i\phi}a,$ (7)
@@
$U^{\dagger}(\phi,2)aU(\phi,2)=e^{i\phi}e^{i\phi(2a^{\dagger}a)}a.$ (8)
The transformation relation for $U(\phi,2)$ in Eq. (8) can be derived using methods from Ref. 40 based on Fock-state representation, with detailed steps provided in Appendix A.
@@
$B_{f}^{\dagger}\begin{pmatrix}a\\ a_{f}\end{pmatrix}B_{f}=\begin{pmatrix}\sqrt{1-l}&\sqrt{l}\\ -\sqrt{l}&\sqrt{1-l}\end{pmatrix}\begin{pmatrix}a\\ a_{f}\end{pmatrix}$, (9)
@@
where $a_{f}$ is the photon annihilation operator for the auxiliary mode $f$ containing vacuum noise $\lvert 0\rangle_{f}$, while $B_{f}$ and $l$ represent the equivalent operator and reflectivity of the virtual beam splitter, respectively. The parameter $l$ corresponds to the photon loss rate, satisfying $0\leq l\leq 1$. Higher $l$ values indicate increased loss, with $l=0$ representing the lossless case and $l=1$ corresponding to complete absorption.
@@
Within this KMZI framework, intensity-difference detection is implemented by placing photon counters at the output ports. Since intensity-difference detection is relatively straightforward to implement experimentally, it finds widespread use for phase sensitivity measurements in MZIs.
III Phase Sensitivity of Intensity-Difference Detection
@@
The phase sensitivity for intensity-difference detection can be expressed using the error propagation formula:
@@
$\Delta\phi=\frac{\sqrt{\langle I_{D}^{2}\rangle-\langle I_{D}\rangle^{2}}}{\bigl\lvert\frac{\partial\langle I_{D}\rangle}{\partial\phi}\bigr\rvert}$, (10)
@@
where $I_{D}=a^{\dagger}a-b^{\dagger}b$, and $\langle\cdot\rangle=\langle\text{out}\rvert\cdot\lvert\text{out}\rangle$ denotes the expectation value with respect to the output state. From the phase estimation model description, the output state can be expressed as:
$\lvert\text{out}\rangle=B_{2}U(\phi,k)B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}.$ (11)
@@
We focus on phase sensitivity for both $k=1$ and $k=2$ cases. For brevity, explicit expressions for phase sensitivity are omitted here. Substituting the expectation values $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ into Eq. (10) yields the phase sensitivity. The specific expressions for $\langle I_{D}\rangle$ and $\langle I_{D}^{2}\rangle$ under $k=1$ and $k=2$ conditions are provided in Appendices B and C, respectively.
@@
Figure 2: Phase sensitivity $\Delta\phi$ versus phase shift $\phi$ for squeezing parameter $r=0.9$ and coherent amplitude $\alpha=1$. (a) and (b): linear phase shift ($k=1$); (c) and (d): Kerr nonlinear phase shift ($k=2$). The black solid line corresponds to $m=n=0$ (no photon addition).
@@
Fig. 2 illustrates the variation of phase sensitivity $\Delta\phi$ with phase shift $\phi$ for $k=1$ and $k=2$ cases, where Figs. 2(a) and (b) depict linear phase shift ($k=1$), while Figs. 2(c) and (d) represent Kerr nonlinear phase shift ($k=2$). According to our computational expressions and the behavior of $\Delta\phi$ in Fig. 2(a), the variation period is $\pi$. Comparing Figs. 2(a) and (c), curves incorporating photon-addition operations show significant improvement over the black solid line ($m=n=0$). Thus, for schemes employing PA-TMSCS input states, photon addition effectively enhances phase sensitivity. Notably, the phase sensitivity variation range in Fig. 2(c) is substantially reduced compared to Fig. 2(a) due to the Kerr nonlinear effect at $k=2$ (hence Fig. 2(c) is not plotted over a full period), suggesting that Kerr nonlinear phase shifting can significantly improve phase sensitivity within certain phase shift ranges.
@@
To further investigate effective photon-addition schemes, we analyze phase sensitivity characteristics under narrowed plotting ranges, as shown in Figs. 2(b) and (d). Phase sensitivity improves with increasing photon-addition order $m+n$, and near the optimal phase sensitivity point, symmetric photon addition ($m=n$) yields the best performance for a given $m+n$. This indicates that performing symmetric photon-addition operations in both $a$- and $b$-modes is more conducive to enhancing measurement accuracy. Consequently, we focus on the $m=n$ scenario for photon-addition operations throughout this paper.
@@
Figure 3: For fixed phase shift $\phi=3.12$: (a) phase sensitivity $\Delta\phi$ versus squeezing parameter $r$ with coherent amplitude $\alpha=1$; (b) $\Delta\phi$ versus $\alpha$ with $r=0.9$. Solid (dashed) lines correspond to $k=2$ ($k=1$).
@@
Fig. 3 demonstrates that phase sensitivity $\Delta\phi$ improves with increasing squeezing parameter $r$, coherent amplitude $\alpha$, and photon-addition order. Additionally, comparing solid and dashed lines clearly shows that Kerr nonlinear phase shift provides improved phase sensitivity compared to linear phase shift.
@@
In practical scenarios, photon loss constitutes an inevitable factor. We subsequently analyze its impact on phase sensitivity and compare achievable measurement precision limits. The measurement precision limit can be defined through the average photon number inside the interferometer $\bar{N}$, expressed as:
$\bar{N}$ $=$ ${}_{\text{int}}\langle\psi\rvert a^{\dagger}a+b^{\dagger}b\lvert\psi\rangle_{\text{int}}$ (12)
$=$ $D_{1}(1,0,1,0)+D_{1}(0,1,0,1),$
@@
where $\lvert\psi\rangle_{\text{int}}=U(\phi,k)B_{f}B_{1}\lvert\text{in}\rangle\lvert 0\rangle_{f}$, and specific expressions for $D_{1}(1,0,1,0)$ and $D_{1}(0,1,0,1)$ are given in Eq. (B5) of Appendix B.
@@
Figure 4: Linear phase shift case ($k=1$): phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL and HL are shown for comparison.
@@
Figure 5: Kerr nonlinear phase shift case ($k=2$): phase sensitivity $\Delta\phi$ versus loss rate $l$ for (a) $m=n=0$, (b) $m=n=1$, (c) $m=n=2$, with fixed phase shift $\phi=3.12$, squeezing parameter $r=0.9$, and coherent amplitude $\alpha=1$. SQL, HL, sub-HL, and SHL are shown for comparison.